# Gas-phase metallicity gradients in galaxies at Key Words.: Galaxies: high-redshift ā Galaxies: abundances ā Galaxies: ISM ā Galaxies: evolution ā Techniques: imaging spectroscopy ā Techniques: high angular resolution1 Introduction

### v1

@@ 2.2.2 SED fitting
With the goal of obtaining the stellar mass of each target, we fitted the spatially integrated spectra extracted from the data cube by using the spectral energy distribution (SED) fitting code bagpipes (Carnall etĀ al., 2018). We adopted the stellar population models by Bruzual & Charlot (2003) and included the nebular emission with cloudy (Ferland etĀ al., 2017) with the ionisation parameter () as a free parameter and adopting a . We assumed a Kroupa (2001) initial mass function truncated at 0.01 and 100 M and a Calzetti etĀ al. (2000) attenuation curve. Finally, we used a non-parametric star-formation history model with continuity priors (see Leja etĀ al. 2019) and with four time bins: Myr, Myr, Myr, and Myr.
@@ 3 A2744-YD4
The top-left panel of Fig.Ā 2 shows the map of the continuum emission in the range 2ā3 m observed wavelength (0.23ā0.34Ā m rest-frame). In addition to the sources named YD4, YD6, YD1, s1, YD7-E, and ZD1, the continuum map also shows the presence of a source at the northernmost part of the FOV, which is also present in the NIRCam image but is not labelled in Hashimoto etĀ al. (2023); we name this source s2. However, given that this is at the edge of the FOV, where many artefacts are present in the NIRSpec IFU data, we consider this source as tentative and report its label in parenthesis in the figure. Moreover, in the map of the continuum obtained from the unsmoothed data cube integrated over the spectral range 1.2ā2 m (Fig. 10, top-left), where the spatial resolution is the highest, the eastern tail of YD1 appears as a separate peaked spatial component. We label this extra source as YD1-E.
@@
Fig. 2, bottom panel, illustrates the integrated spectrum (aperture radius = 0.15) associated with the main component of the A2744-YD4 system, that is, YD4 itself. The rest-frame optical lines [Oāii], H, and [Oāiii] are detected with high S/N (6 on the peak). The spectroscopic redshift based on Lyman break and optical lines ( 7.88) is consistent with that found by Morishita etĀ al. (2023) and Hashimoto etĀ al. (2023) from the NIRSpec 2700 data. We find a stellar mass of log() 8.7 from the SED fitting and a SFR 2 Ā yr from H for YD4 (Table 3). These were corrected for the magnification factor of 2 due to the lensing. The gas-phase metallicity, relying on line ratios, is instead not affected by it. For the other spatial components in the system, we find log() 7.5ā8.6 and SFR 1ā3 Ā yr.
@@
The [Oāiii] flux map resulting from our spectral emission-line modelling is displayed in Fig. 2, central panel (the maps of H and [Oāii] are shown in Fig. 10). From the [Oāiii] map, we identify two new emission regions which are weak or absent in continuum, which we label as YD6-[OIII]-E and YD4-[OIII]-W. Overall, the [Oāiii]Ā maps obtained from the PRISM data presented here are deeper and reveal fainter and more extended features than those from high-resolution grating data presented in Hashimoto etĀ al. (2023) (who do not report the detection of any other emission line due to the lower S/N of their data). We can count nine sources in total in the 3 NIRSpec IFU FOV, among those detected in continuum and in line emission, namely YD4, YD6, YD6-[OIII]-E, YD4-[OIII]-W, YD7-E, ZD1, YD1, YD1-E, and s1, and possibly a tenth source, s2. All these sources lie at 7.88, either confirmed spectroscopically (YD4, YD6, YD6-[OIII]-E, YD4-[OIII]-W, YD1, YD1-E, and s1; e.g. Morishita etĀ al. 2023, Hashimoto etĀ al. 2023, and this work) or from photometry (ZD1), except for YD7-E (and the tentative source s2) whose redshift is not assessed. The spectra of all targets, but YD4, are reported in Figs.Ā 8 and 9.
@@
In the top-right panel of Fig. 2 we show the map of oxygen abundance, 12 + ā(O/H), a proxy for gas-phase metallicity. The metallicity is inferred by making use of the strong-line diagnostics reported in Table 2 as described in Sect. 2.2 (the maps of the [Oāiii] 5008/HĀ and [Oāiii] 5008/[Oāii] 3727,30Ā emission line ratios are shown in Fig. 10). The ratios involving lines which are not close in wavelength (all but [Oāiii]Ā and H) may be affected by extinction. Unfortunately, the only robustly detected Balmer line at the spaxel level is H, and only in a handful of spaxels, therefore estimating the dust extinction spaxel-by-spaxel is not possible. Based on this, the abundance map should not be taken as a robust measurement; moreover, the map is very noisy. Nevertheless, the map seems to suggest that, in the northern system, YD1 is embedded between two sources at lower metallicity, s1 and YD1-E. No clear pattern of metallicity shows up in the southern system (the YD4 one) from the map.
@@ 4 BDF-3299
BDF-3299 is a spectroscopically confirmed star-forming galaxy (SFR 6 Ā yr, SFR 12 /yr) at 7.109 (Castellano etĀ al., 2010; Vanzella etĀ al., 2011; Maiolino etĀ al., 2015; Carniani etĀ al., 2017), located in an over-density of galaxies (Castellano etĀ al., 2016). We report the maps for BDF-3299 from the NIRSpec IFU observations in Fig. 3.
@@
Fig. 3, bottom panel, shows the integrated spectrum (aperture radius = 0.1) associated with the brightest component in continuum emission, BDF-3299-a. The rest-frame optical lines [Oāii], H, and [Oāiii] are detected with high S/N (6 on the peak). From the lines, we estimate a redshift of 7.114, slightly higher but roughly consistent with the previously reported one from Ly (7.109; Vanzella etĀ al., 2011) and [Cāii] (7.107; Carniani etĀ al., 2017). We obtain a stellar mass of log() 7.9 from the SED fitting and a SFR 3.5 Ā yr from H for BDF-3299-a (Table 3). The spectra of the other spatial components in the system, namely BDF-3299-b and c, are reported in Fig.Ā 11. From these, we find log() 7.6ā8.2 and SFR 1.5ā4 Ā yr.
The [Oāiii] emission (Fig. 3, central panel) peaks on BDF-3299-b rather than on BDF-3299-a (the strongest in continuum emission). The line ratio maps are quite noisy and only few spaxels are above the S/N threshold of 3 (Fig. 12). Therefore, the same applies to the metallicity map, which generally shows low values, of 12+log(O/H) 7.8, in the few spaxels where it can be estimated (Fig. 3, right).
@@ 5 COSMOS-24108
COSMOS-24108 (Ā Ā 6.36; SFR 29 /yr, SFR 6.2 /yr) shows two, or possibly three, spatial components in H-band HST rest-frame UV imaging (Pentericci etĀ al., 2016). In Fig. 4, left panel, we show the 1ā2 m observed continuum (0.14ā0.27 m rest-frame) from the NIRSpec IFU data. This shows two main spatial components, and a third, fainter one in the SE part of the system, consistent with those seen with HST in a similar spectral band. We label these three spatial components as COSMOS24108-a, b, and c, in order of continuum brightness. The continuum at redder wavelengths, between 3ā3.5 m (0.39ā0.46 m rest-frame; Fig. 14), is dominated by the southernmost of the two main components, and only an extended emission towards the northernmost source is present at these wavelengths, instead of a more clearly separate spatial component as at lower wavelengths. Therefore, the northernmost source has bluer continuum than the southernmost one. In the SW corner of the FOV, another galaxy at a lower redshift is also present.
@@
Fig. 4, bottom panel, shows the integrated spectrum (aperture radius = 0.15) associated with the brightest component in continuum emission, COSMOS24108-a. The rest-frame optical lines [Oāii], H, [Oāiii], H, and [Sāii] are detected with high S/N (4 on the peak). We obtain a stellar mass of log() 9.3 from SED fitting and a SFR 15 Ā yr from H for COSMOS24108-a (Table 3). The integrated spectra of the other spatial components in the system are shown in Fig.Ā 13. For these, we find log() 8.4ā8.9 and SFR 2ā15 Ā yr.
@@
The [Oāiii] ionised gas line emission (Fig. 4, central panel; H, [Oāii], and H maps are reported in Fig. 14) is much more extended than the continuum, revealing two additional bright clumps to the SE of the system, one of the two also tentatively detected in continuum (see contours). We label these as COSMOS24108-[OIII]-Ea and [OIII]-Eb. The HĀ map (Fig. 14, bottom-left) shows a bridge of gas (visible also in [Oāiii], though weaker relative the rest of the emission) connecting the main system to the closer of these two clumps, COSMOS24108-[OIII]-Ea. By comparing the ionised gas line emission with the 1ā2 m continuum, we see that the former peaks in between the two main continuum components, with a preferential extension towards the northernmost, weaker continuum component (COSMOS24108-b) rather than to the southernmost, brighter one (COSMOS24108-a).
@@
Fig.Ā 4, right panel, reports the map of gas metallicity (the [Oāiii] 5008/H and [Oāiii] 5008/[Oāii] 3727,30 line ratio maps are shown in Fig. 14). The metallicity exhibits differences of up to 0.5ā0.6 dex among the different sources in the system. The northernmost of the two main continuum sources, COSMOS24108-b, appears to have lower metallicity (12 + ā(O/H) 8.0ā8.1) than the southernmost one, COSMOS24108-a ( 8.2), while the SE minor continuum component, COSMOS24108-c, has larger metallicity ( 8.4). Towards the two ionised gas clumps to the SE, COSMOS24108-[OIII]-Ea and COSMOS24108-[OIII]-Eb, the metallicity is the lowest, with values of 7.8ā7.9.
@@ 6 Integrated mass-metallicity relation
We obtained the properties of each sub-source belonging to each target, by extracting integrated spectra from circular apertures of radius of 0.15 for A2744-YD4 and COSMOS24108 and of 0.1 for BDF-3299 centred on each sub-source (Table 3). We find emission line ratios of log([Oāiii]/H) 0.4ā0.9 for A2744-YD4, log([Oāiii]/H) 0.7ā0.8 for BDF-3299, and log([Oāiii]/H) 0.5ā0.9 and log([Sāii]/H) ā0.8 for COSMOS24108. These are consistent with the range of values found for 6ā8 star-forming galaxies in the JADES survey (Cameron etĀ al., 2023), of log([Oāiii]/H) 0.5ā0.8 and log([Sāii]/H) ā0.8 (when only considering detections and not upper limits), in a range of stellar masses (log() 6.5ā9.0) and SFRs (0.1ā30) comprising those of the sources in this work. In some cases, like COSMOS24108-a with log([Oāiii]/H) 0.8 and log([Sāii]/H) ā0.8, the ratios are at the high end of the values reported in (Cameron etĀ al., 2023).
@@
We inferred the gas-phase metallicity and the SFR from emission-line fitting and the stellar mass from SED fitting, as described in Sect. 2.2. In Fig.Ā 5 we display the metallicity versus stellar mass diagram for the spatial components identified in the images (TableĀ 3). We compare these with the values reported by Curti etĀ al. (2023b) for the 3ā10 galaxies from the JADES (including GNz-11; Bunker etĀ al. 2023), CEERS (Nakajima etĀ al., 2023, re-computed by Curti etĀ al. 2023b for consistency), and EROs (Curti etĀ al., 2023a; Laseter etĀ al., 2024) samples (grey circles). The best-fit relation for this 3ā10 compilation obtained by Curti etĀ al. (2023b) is also shown. We also report for reference the best-fit mass-metallicity relations (MZRs) for galaxies at 0.08 from SDSS (Curti etĀ al., 2020b), 2ā3 from MOSDEF (Sanders etĀ al., 2021), and 3 from GLASS (Li etĀ al., 2023).
@@
Our targets are generally compatible with the values found for the 3ā10 galaxies, especially the sources in BDF-3299 which sit on the MZR defined by this high- collection. We note that most of the sources in A2744-YD4 and COSMOS24108 ( 8 and 6, respectively) are at the high end of the JADES, CEERS, and EROs points, where the 2ā3 MZRs from MOSDEF and GLASS lie. This suggests that the A2744-YD4 and COSMOS24108 systems may comprise more evolved sources as compared to the majority of sources at the same redshifts and be instead more similar to galaxies at cosmic noon.
@@
We further display the predictions for the high- MZR from a number of cosmological simulations, specifically EAGLE ( 8; Schaye etĀ al., 2015), FIRE ( 6; Ma etĀ al., 2016), IllustrisTNG ( 8; Torrey etĀ al., 2019), FirstLight ( 6 and 8; Langan etĀ al., 2020), and SERRA ( 8; Pallottini etĀ al., 2022)111The MZR from the SERRA simulation (Pallottini etĀ al., 2022) was originally presented in Curti etĀ al. (2023a) and was computed using 202 simulated galaxies at . In the present work, the MZR from SERRA is recomputed from an extended sample of 245 objects Pallottini & Ferrara (2023). This causes some differences at the low-mass end of the reported MZR with respect to that in Curti etĀ al. (2023a).. We note that the sources analysed in this work, both as a whole and within the individual A2744-YD4 and COSMOS24108 systems, are better aligned with simulations predicting steeper MZR slopes than the median of the observations at 3ā10.

### v2

@@ 2.2.2 SED fitting
With the goal of obtaining the stellar mass of each target, we fitted the spatially integrated spectra extracted from the data cube by using the SED fitting code bagpipes (Carnall etĀ al., 2018). We adopted the stellar population models by Bruzual & Charlot (2003) and included the nebular emission with cloudy (Ferland etĀ al., 2017) with the ionisation parameter () as a free parameter and adopting a . We assumed a Kroupa (2001) initial mass function truncated at 0.01 and 100 Mā and a Calzetti etĀ al. (2000) attenuation curve. Finally, we used a non-parametric star-formation history model with continuity priors (see Leja etĀ al. 2019) and with four time bins: Myr, Myr, Myr, and Myr.
@@ 3 A2744-YD4
The top-left panel of Fig.Ā 2 shows the map of the continuum emission in the range 2ā3 m observed wavelength (0.23ā0.34Ā m rest-frame). In addition to the sources named YD4, YD6, YD1, s1, YD7-E, and ZD1, the continuum map also shows the presence of a source at the northernmost part of the FOV, which is also present in the NIRCam image but is not labelled in Hashimoto etĀ al. (2023); we name this source s2. However, given that this is at the edge of the FOV, where many artefacts are present in the NIRSpec IFU data, we consider this source as tentative and report its label in parenthesis in the figure. Moreover, in the map of the continuum obtained from the unsmoothed data cube integrated over the spectral range 1.2ā2 m (Fig.Ā A in the Appendix333The Appendix is available at https://doi.org/10.5281/zenodo.13327942., top-left), where the spatial resolution is the highest, the eastern tail of YD1 appears as a separate peaked spatial component. We label this extra source as YD1-E.
@@
Fig. 2, bottom panel, illustrates the integrated spectrum (aperture radius = 0.15ā²ā²) associated with the main component of the A2744-YD4 system, that is, YD4 itself. The rest-frame optical lines [Oāii], H, and [Oāiii] are detected with high S/N (6 on the peak). The spectroscopic redshift based on Lyman break and optical lines ( 7.88) is consistent with that found by Morishita etĀ al. (2023) and Hashimoto etĀ al. (2023) from the NIRSpec 2700 data. We find a stellar mass of log() 8.7 from the SED fitting and a SFR 2 Ā yr-1 from H for YD4 (Table 2). These were corrected for the magnification factor of 2 due to the lensing. The gas-phase metallicity, relying on line ratios, is instead not affected by it. For the other spatial components in the system, we find log() 7.5ā8.6 and SFR 1ā3 Ā yr-1.
@@
The [Oāiii] flux map resulting from our spectral emission-line modelling is displayed in Fig. 2, central panel (the maps of H and [Oāii] are shown in Fig.Ā Aā ā footnotemark: ). From the [Oāiii] map, we identify two new emission regions which are weak or absent in continuum, which we label as YD6-[OIII]-E and YD4-[OIII]-W. Overall, the [Oāiii]Ā maps obtained from the PRISM data presented here are deeper and reveal fainter and more extended features than those from high-resolution grating data presented in Hashimoto etĀ al. (2023) (who do not report the detection of any other emission line due to the lower S/N of their data). We can count nine sources in total in the 3 NIRSpec IFU FOV, among those detected in continuum and in line emission, namely YD4, YD6, YD6-[OIII]-E, YD4-[OIII]-W, YD7-E, ZD1, YD1, YD1-E, and s1, and possibly a tenth source, s2. All these sources lie at 7.88, either confirmed spectroscopically (YD4, YD6, YD6-[OIII]-E, YD4-[OIII]-W, YD1, YD1-E, and s1; e.g. Morishita etĀ al. 2023, Hashimoto etĀ al. 2023, and this work) or from photometry (ZD1), except for YD7-E (and the tentative source s2) whose redshift is not assessed. The spectra of all targets, but YD4, are reported in Figs.Ā A and A in the Appendix444The Appendix is available at https://doi.org/10.5281/zenodo.13327942..
@@
In the top-right panel of Fig. 2, we show the map of oxygen abundance, 12 + ā(O/H), a proxy for gas-phase metallicity. The metallicity is inferred by making use of the strong-line diagnostics reported in Table 2 as described in Sect. 2.2 (the maps of the [Oāiii] 5008/HĀ and [Oāiii] 5008/[Oāii] 3727,30Ā emission line ratios are shown in Fig. Aā ā footnotemark: ). The ratios involving lines which are not close in wavelength (all but [Oāiii]Ā and H) may be affected by extinction. Unfortunately, the only robustly detected Balmer line at the spaxel level is H, and only in a handful of spaxels, therefore estimating the dust extinction spaxel-by-spaxel is not possible. Based on this, the abundance map should not be taken as a robust measurement; moreover, the map is very noisy. Nevertheless, the map seems to suggest that, in the northern system, YD1 is embedded between two sources at lower metallicity, s1 and YD1-E. No clear pattern of metallicity shows up in the southern system (the YD4 one) from the map.
@@ 4 BDF-3299
BDF-3299 is a spectroscopically confirmed star-forming galaxy (SFRUV 6 Ā yr-1, SFRdust 12 /yr) at 7.109 (Castellano etĀ al., 2010; Vanzella etĀ al., 2011; Maiolino etĀ al., 2015; Carniani etĀ al., 2017), located in an over-density of galaxies (Castellano etĀ al., 2016). We report the maps for BDF-3299 from the NIRSpec IFU observations in Fig. 3.
@@
Fig. 3, bottom panel, shows the integrated spectrum (aperture radius = 0.1ā²ā²) associated with the brightest component in continuum emission, BDF-3299-a. The rest-frame optical lines [Oāii], H, and [Oāiii] are detected with high S/N (6 on the peak). From the lines, we estimate a redshift of 7.114, slightly higher but roughly consistent with the previously reported one from Ly (7.109; Vanzella etĀ al., 2011) and [Cāii] (7.107; Carniani etĀ al., 2017). We obtain a stellar mass of log() 7.9 from the SED fitting and a SFR 3.5 Ā yr-1 from H for BDF-3299-a (Table 2). The spectra of the other spatial components in the system, namely BDF-3299-b and c, are reported in Fig.Ā B in the Appendixā ā footnotemark: ). From these, we find log() 7.6ā8.2 and SFR 1.5ā4 Ā yr-1.
The [Oāiii] emission (Fig. 3, central panel) peaks on BDF-3299-b rather than on BDF-3299-a (the strongest in continuum emission). The line ratio maps are quite noisy and only few spaxels are above the S/N threshold of 3 (Fig.Ā B). Therefore, the same applies to the metallicity map, which generally shows low values, of 12+log(O/H) 7.8, in the few spaxels where it can be estimated (Fig. 3, right).
@@ 5 COSMOS-24108
COSMOS-24108 (Ā Ā 6.36; SFRUV 29 /yr, SFRdust 6.2 /yr) shows two, or possibly three, spatial components in H-band HST rest-frame UV imaging (Pentericci etĀ al., 2016). In Fig.Ā 4, left panel, we show the 1ā2 m observed continuum (0.14ā0.27 m rest-frame) from the NIRSpec IFU data. This shows two main spatial components, and a third, fainter one in the SE part of the system, consistent with those seen with HST in a similar spectral band. We label these three spatial components as COSMOS24108-a, b, and c, in order of continuum brightness. The continuum at redder wavelengths, between 3ā3.5 m (0.39ā0.46 m rest-frame; Fig.Ā C in the Appendixā ā footnotemark: ), is dominated by the southernmost of the two main components, and only an extended emission towards the northernmost source is present at these wavelengths, instead of a more clearly separate spatial component as at lower wavelengths. Therefore, the northernmost source has bluer continuum than the southernmost one. In the SW corner of the FOV, another galaxy at a lower redshift is also present.
@@
Fig. 4, bottom panel, shows the integrated spectrum (aperture radius = 0.15ā²ā²) associated with the brightest component in continuum emission, COSMOS24108-a. The rest-frame optical lines [Oāii], H, [Oāiii], H, and [Sāii] are detected with high S/N (4 on the peak). We obtain a stellar mass of log() 9.3 from SED fitting and a SFR 15 Ā yr-1 from H for COSMOS24108-a (Table 2). The integrated spectra of the other spatial components in the system are shown in Fig.Ā C in the Appendixā ā footnotemark: . For these, we find log() 8.4ā8.9 and SFR 2ā15 Ā yr-1.
@@
The [Oāiii] ionised gas line emission (Fig. 4, central panel; H, [Oāii], and H maps are reported in Fig. Cā ā footnotemark: ) is much more extended than the continuum, revealing two additional bright clumps to the SE of the system, one of the two also tentatively detected in continuum (see contours). We label these as COSMOS24108-[OIII]-Ea and [OIII]-Eb. The HĀ map (Fig.Ā C, bottom-left) shows a bridge of gas (visible also in [Oāiii], though weaker relative the rest of the emission) connecting the main system to the closer of these two clumps, COSMOS24108-[OIII]-Ea. By comparing the ionised gas line emission with the 1ā2 m continuum, we see that the former peaks in between the two main continuum components, with a preferential extension towards the northernmost, weaker continuum component (COSMOS24108-b) rather than to the southernmost, brighter one (COSMOS24108-a).
@@
Fig.Ā 4, right panel, reports the map of gas metallicity (the [Oāiii] 5008/H and [Oāiii] 5008/[Oāii] 3727,30 line ratio maps are shown in Fig. Cā ā footnotemark: ). The metallicity exhibits differences of up to 0.5ā0.6 dex among the different sources in the system. The northernmost of the two main continuum sources, COSMOS24108-b, appears to have lower metallicity (12 + ā(O/H) 8.0ā8.1) than the southernmost one, COSMOS24108-a ( 8.2), while the SE minor continuum component, COSMOS24108-c, has larger metallicity ( 8.4). Towards the two ionised gas clumps to the SE, COSMOS24108-[OIII]-Ea and COSMOS24108-[OIII]-Eb, the metallicity is the lowest, with values of 7.8ā7.9.
@@ 6 Integrated mass-metallicity relation
We obtained the properties of each sub-source belonging to each target, by extracting integrated spectra from circular apertures of radius of 0.15ā²ā² for A2744-YD4 and COSMOS24108 and of 0.1ā²ā² for BDF-3299 centred on each sub-source (Table 2). We find emission line ratios of log([Oāiii]/H) 0.4ā0.9 for A2744-YD4, log([Oāiii]/H) 0.7ā0.8 for BDF-3299, and log([Oāiii]/H) 0.5ā0.9 and log([Sāii]/H) ā0.8 for COSMOS24108. These are consistent with the range of values found for 6ā8 star-forming galaxies in the JADES survey (Cameron etĀ al., 2023), of log([Oāiii]/H) 0.5ā0.8 and log([Sāii]/H) ā0.8 (when only considering detections and not upper limits), in a range of stellar masses (log() 6.5ā9.0) and SFRs (0.1ā30) comprising those of the sources in this work. In some cases, like COSMOS24108-a with log([Oāiii]/H) 0.8 and log([Sāii]/H) ā0.8, the ratios are at the high end of the values reported in (Cameron etĀ al., 2023).
@@
We inferred the gas-phase metallicity and the SFR from emission-line fitting and the stellar mass from SED fitting, as described in Sect. 2.2. In Fig.Ā 5, we display the metallicity versus SFR and stellar mass diagrams for the spatial components identified in the images (TableĀ 2). We compare these with the values reported by Curti etĀ al. (2023b) for the 3ā10 galaxies from the JADES (including GNz-11; Bunker etĀ al. 2023), CEERS (Nakajima etĀ al., 2023, re-computed by Curti etĀ al. 2023b for consistency), and EROs (Curti etĀ al., 2023a; Laseter etĀ al., 2024) samples (grey circles). The best-fit relation for this 3ā10 compilation obtained by Curti etĀ al. (2023b) is also shown. We also report for reference the best-fit mass-metallicity relations (MZRs) for galaxies at 0.08 from SDSS (Curti etĀ al., 2020b), 2ā3 from MOSDEF (Sanders etĀ al., 2021), and 3 from GLASS-JWST (Li etĀ al., 2023).
@@
Our targets are generally compatible with the values found for the 3ā10 galaxies, especially the sources in BDF-3299 which sit on the MZR defined by this high- collection. We note that most of the sources in A2744-YD4 and COSMOS24108 ( 8 and 6, respectively) are at the high end of the JADES, CEERS, and EROs points, where the 2ā3 MZRs from MOSDEF and GLASS-JWST lie. This suggests that the A2744-YD4 and COSMOS24108 systems may comprise more evolved sources as compared to the majority of sources at the same redshifts and be instead more similar to galaxies at cosmic noon.
@@
We further display the predictions for the high- MZR from a number of cosmological simulations, specifically EAGLE ( 8; Schaye etĀ al., 2015), FIRE ( 6; Ma etĀ al., 2016), IllustrisTNG ( 8; Torrey etĀ al., 2019), FirstLight (from Langan etĀ al. 2020 at 6 and 8 and Nakazato etĀ al. 2023 at 8, for 9 and 9, respectively, at 6), and SERRA ( 8; Pallottini etĀ al., 2022)555The MZR from the SERRA simulation (Pallottini etĀ al., 2022) was originally presented in Curti etĀ al. (2023a) and was computed using 202 simulated galaxies at . In the present work, the MZR from SERRA is recomputed from an extended sample of 245 objects Pallottini & Ferrara (2023). This causes some differences at the low-mass end of the reported MZR with respect to that in Curti etĀ al. (2023a).. We note that the sources analysed in this work, both as a whole and within the individual A2744-YD4 and COSMOS24108 systems, are better aligned with simulations predicting steeper MZR slopes than the median of the observations at 3ā10.